What does y= in the slope intercept form of the linear equation containing the given pair of points (-1,-3) and (-9,-5)
Answer by rfer(16322)

y=mx+b
Slope
m=(-5--3)/(-9--1)
m=-2/-8
m=1/4
-------------
-5=1/4*-9+b
-5=-9/4+b
b=9/4-5/1
b=9/4-20/4
b=-11/4
b=-2 3/4
-----------
y=1/4x-2 3/4